A d, c motor consumes 47.25MJ when connected to a 250 V supply for 1 hour 45 minutes. Determine the power rating of the motor and the current taken from the supply

1
Expert's answer
2022-03-21T12:45:37-0400

We know that

Power

"P=\frac{W}{t}"

"P=\frac{47.25\times10^6}{6300}=7500W"

Now we know that

"P=iV"

"I=\frac{P}{V}=\frac{7500}{250}=30A"